4. A composite floor system consists of a 5-inch slab which is supported by W14x30 beams without shoring. Fy-50 ksi, strength of concrete is 4ksi, beam spacing is 8-ft, the effective width is 81-inches, and the span is 35-ft. Construction load is 20 psf and the live load is 160 psf. Compute the maximum deflection before the concrete has cured. (credit weight 20)

Please refer AISC 15th edition
Answer #1

since the deflection required is before curing, there is not composite action and the loads are only self weight of slab and construction load

given beam is W14x30

slab thickness = 5 in

self weight of slab = (5/12)*150=62.5 psf

beam spacing = 8 ft

construction live load = 20 psf

tributary width = 8 ft

dead load on beam from slab = 62.5*8=500 lb/ft = 0.5 kip/ft

live load on beam = 20*8=160 lb/ft = 0.16 kip/ft

self weight of beam = 30 lb/ft = 0.03 kip/ft

total load on beam = 0.5+0.16+0.03 = 0.69 kip/ft = 0.0575 kip/in

span of beam = 35 ft= 420 in

moment of inertia of beam = 291 in4

deflection of beam = (5/384)*0.0575*4204/29000*291=2.76 inches

2.76 inches is maximum deflection of beam before curing